# Decoding Intelligence: A Framework for Certifying Knowledge Comprehension in LLMs

### v2

@@ B.3 Prompt Construction
These inclusions ensure that the prompt is comprehensive, fits within model constraints, and provides sufficient guidance for the LLM to generate accurate responses.
B.3.1 Distractor Selection
Distractors are crucial in assessing an LLM’s comprehension and reasoning abilities. We hypothesize that distractors to nodes later in the path closer to the answer would provide more difficulty for the LLM due to their proximity to the answer node. Our distractor selection process, implemented in Algorithm 4, employs a weighted sampling approach to prioritize distractors associated with entities closer to the path’s end. The algorithm first identifies all potential distractors for each node in the path, then assigns weights inversely proportional to their distance from the final answer node. This weighting scheme favors distractors linked to nodes near the path’s end, but still allows for the selection of distractors related to any node in the path. This refers to the sampling procedure in the line 12 of the probabilistic progam specification  1. By performing weighted sampling from this pool, we ensure a balance between highly relevant distractors and a diverse selection across the entire path.
B.3.2 Answer Options
@@
The process of generating answer options is detailed in Algorithm 3. In the algorithm, we sample answer options from the set described above so we are basically sampling from the nodes as in the probabilistic program specifiction line 12  1. The answer option algorithm assumes that distractors are input in a list according to the order of preference.
@@
1:  Input: Graph $G$, Path $\Pi$
@@
5:  for $i\leftarrow$ 00 to len($\Pi$) $-2$ do
6:     $v\leftarrow\Pi[i]$
@@
8:     $N\_distractors\leftarrow$ FilterDistractors($N$, $v$, $\Pi$)
@@ B.4 Context Trimming
B.4.1 Finding Relevant Sentences per Edge
@@
B.4.2 Trimming to Fit Context Length
@@
B.5 Few-Shot Examples
To guide the LLM towards the desired response format and demonstrate the reasoning process, we include 2 few-shot examples in the prompt. These examples provide a clear illustration of how to approach the multi-hop reasoning task.
We use the following few-shot examples:
@@ B.5 Few-Shot Examples
B.6 Final Prompt
The final prompt presented to the LLM is constructed using a template that incorporates several key elements:
Trimmed Context [B.4]: The relevant context extracted and trimmed.
Query [B.2]: The multi-hop query.
Answer Options [B.3.2]: The generated answer options, including the correct answer and distractors.
Few-Shot Examples [B.5]: A set of examples demonstrating the desired response format.
@@ B.6 Final Prompt
B.7 Response Checker function

### v3

@@ D.5 Prompt Construction
These inclusions ensure that the prompt is comprehensive, fits within model constraints, and provides sufficient guidance for the LLM to generate accurate responses. We also provide information on the aliases used and the entities they correspond to in the prompt, to ensure that the LLM knows about the alias.
D.5.1 Distractor Selection
Following Definition 2.2, we implement distinct distractor selection strategies for each knowledge graph structure to create challenging evaluation scenarios.
Wikidata5m Linear Paths
For a path ${\mathcal{N}}=[v_{1},...,v_{k}]$, we select distractors through weighted sampling based on path position:
1. Distractor Identification: For each node $v_{i}\in{\mathcal{N}}$: - Find candidate distractors $D_{i}=\{\tilde{v}\mid(v_{i},\tilde{v})\in{\mathcal{E}}\land\tilde{v}\notin{\mathcal{N}}\}$ Filter to preserve relation type: $D_{i}^{\prime}=\{\tilde{v}\in D_{i}\mid\exists j>i:(v_{i},\tilde{v})_{\mathcal{A}}=(v_{i},v_{j})_{\mathcal{A}}\}$
2. Position-based Weighting: Assign weights to $\tilde{v}\in D_{i}^{\prime}$:
$w(\tilde{v})=i/|{\mathcal{N}}|$
This weights distractors near path end higher, as formalized in Algorithm 6.
PrimeKG DAG Structure
For reference DAG ${\mathcal{G}}_{ref}$ and its isomorphic instance ${\mathcal{G}}^{\prime}$, we:
1. Include all valid distractors in context:
[…]
2. Ensure answer options contain at least one distractor:
$\text{options}=\{{\mathcal{N}}^{\prime}_{sink}\}\cup\{d\sim\text{Uniform}(D)\}\cup\text{Others}$
This flexible approach demonstrates how our framework is easily adaptable to different kinds of queries.
D.5.2 Answer Options
@@
The process of generating answer options is detailed in Algorithm 5. In the algorithm, we sample answer options from the set described above. The answer option algorithm assumes that distractors are input in a list according to the order of preference. Another important note is that for primeKG, we try to ensure that all options are of the same type as the correct answer to prevent the LLM from easily eliminating wrong options.
@@
1:  Input: Graph $G$, Path ${\mathcal{N}}$
@@
5:  for $i\leftarrow$ 00 to len(${\mathcal{N}}$) $-2$ do
6:     $v\leftarrow{\mathcal{N}}[i]$
@@
8:     $N\_distractors\leftarrow$ FilterDistractors($N$, $v$, ${\mathcal{N}}$)
@@ D.6 Context Trimming
D.6.1 Finding Relevant Sentences per Edge (Wikidata5m)
@@
D.6.2 Relevant sentences per Edge PrimeKG
As we synthetically generate data for each edge in PrimeKG, we already have a sentence that corresponds to an edge in primeKG. So no further processing is needed.
D.6.3 Trimming to Fit Context Length
@@
Vanilla Setting Note:
In the vanilla setting, we only want to check simple knowledge comprehension abilities of LLMs, so if our DAG used to create a query has an edge type (alias), we ensure that the provided context trims any sentences which are in the relevant sentences of the same edge type has any edge type in the DAG. In distractor setting, there is no such constraint.
@@
D.7 Few-Shot Examples
To guide the LLM towards the desired response format and demonstrate the reasoning process, we include 2 few-shot examples in the prompt (common across all prompts to all models). These examples provide a clear illustration of how to approach the knowledge comprehension task. We investigate the impact of varying number of few-shot examples on LLM performance in Appendix E. We use the following few-shot examples for wikidata5m:
@@ D.7 Few-Shot Examples
We use the following few-shot examples for primeKG:
Few Shot Examples Common Context: (drug) entity_A has side effects: effect_1, effect_5. (disease) entity_C has contraindication relation to (drug) entity_E. (disease) entity_C is indicated for (drug) entity_A. (drug) entity_D has side effects: effect_2, effect_3, effect_4. (drug) entity_D has indication for (disease) entity_C. (drug) entity_E has indication to (disease) entity_B. (gene/protein) entity_F is carrier for (drug) entity_E. (disease) entity_B is indicated for (drug) entity_G. Question 1: Which drug $\rightarrow$ (has indication for) (disease) entity_C and has $\rightarrow$ (side effect) effect_5 ? Options: 1. entity_F, n 2. entity_E, n 3. entity_D, n 4. entity_A, n 5. entity_B Answer: 2. entity_A Explanation: entity_A has the side effects effect_5 while it is the indication of entity_C, so answer is entity_A. How to get answer: Find the drugs entity_C is indicated for, which gives us entity_A and then check which of them has the side effect_5. Question 2: What drugs should be avoided in (disease) entity_C but are indicated for (disease) entity_B? Options: 1. entity_F, n 2. entity_A, n 3. entity_D, n 4. entity_E, n 5. entity_B Answer: 4. entity_E Explanation: entity_B is indicated for entity_E (drug), and entity_E has contraindication to entity_C. How to get answer: find the drugs entity_B is indicated for to get entity_E, and entity_G. then find the contraindications of entity_C to get entity_E. The intersection of these two sets gives the answer, which in this case is entity_E.
D.8 Final Prompt
The final prompt presented to the LLM is constructed using a template (same for wikidata5m and primeKG) that incorporates several key elements:
Trimmed Context [D.6]: The relevant context extracted and trimmed.
Query [D.4]: The multi-hop query.
Answer Options [D.5.2]: The generated answer options, including the correct answer and distractors.
Few-Shot Examples [D.7]: A set of examples demonstrating the desired response format.
@@ D.8 Final Prompt
D.9 Response Checker function
@@
Appendix E Ablations
E.1 Few Shot Prompts
Table 2: Certification results for LLMs in vanilla setting with different number of few-shot examples
Model Avg. lower bound Avg. upper bound Avg. accuracy
Gemini-1.5-Flash 2Shot (Default) $0.46\pm 0.06$ $0.58\pm 0.06$ $0.52\pm 0.06$
Gemini-1.5-Flash 3Shot $0.46\pm 0.06$ $0.58\pm 0.06$ $0.52\pm 0.06$
Gemini-1.5-Flash 4Shot $0.46\pm 0.07$ $0.58\pm 0.07$ $0.52\pm 0.07$
Gemini-1.5-Flash 5Shot $0.46\pm 0.07$ $0.58\pm 0.07$ $0.52\pm 0.07$
E.2 Distractor Distributions
  1. 1.
  2. 2.
  3. 3.
Model Avg. lower bound Avg. upper bound Avg. accuracy
Gemini-1.5-Flash Setting 2 $0.42\pm 0.11$ $0.55\pm 0.11$ $0.48\pm 0.11$
Gemini-1.5-Flash Setting 3 $0.42\pm 0.11$ $0.55\pm 0.11$ $0.48\pm 0.11$
E.3 Model Performances with varying Path Length
Varying Setting: In figure  7 we show plots for various specifications for the GPT4o model.
Varying Quantization: In figure  8 we show plots when the quantization is varied with the Llama3-8B model on the shuffle specification and their effects on performance.
Varying Models: In fig  9 we show plots for the shuffle specification and performance across the models(the open-source models use fp16 precision).
In Figure  7, we observe that the performance across settings converges as k increases and the distractor setting is most impactful on the performance for $k=2$.
In Figure  8, we infer that as k increases the performance of the models’ on the task converges across the different quantizations. We hypothesize this is due to the increasing complexity of the reasoning task.
E.4 Chain of Thought Prompting
We also conduct an ablation on how Chain-of-Thought(COT) prompting can affect the performance of language models on the knowledge comprehension task. Specifically, we investigate the Phi-3 3B model (precision: float16) in the vanilla setting with COT prompting strategy. We augmented our standard few-shot examples (D.7) with COT steps and added structured reasoning guidance to the prompt template (D.5):
COT additions to prompt template Answer in the following the below format: Let’s solve this step by step: 1) Let’s identify the starting point and path: - Start: [identify starting entity] - Path to follow: [break down the path components] 2) Let’s follow the path: Starting from [entity] → [first relationship] → [next entity] → [next relationship] → [next entity] … [continue as needed] 3) Verify our final destination reaches one of the given options Therefore, the correct answer is: $<$option_number$>$. $<$option_text$>$
In the vanilla setting, adding COT prompting improved Phi-3 3B’s performance, with the bounds increasing by 0.11 summarized in Table 4. While we acknowledge the potential benefits of COT, earlier experiments were limited due to the significantly increased computational cost (generating 5-8 times more tokens) and the expenses of COT, particularly with closed-source models as output tokens are much more expensive.
Table 4: Certification results for Phi-3 3B with and without COT
Prompting Strategy Avg. lower bound Avg. upper bound Avg. accuracy
No COT (Default) $0.34\pm 0.05$ $0.46\pm 0.06$ $0.40\pm 0.05$
COT $0.45\pm 0.08$ $0.57\pm 0.08$ $0.51\pm 0.08$
Appendix F Case Studies from certifying over Wikidata5m
We analyze the certification results, qualitatively. First, we show the responses of $3$ models in Figure 10 — Phi-3 (3B), GPT-4o, and Gemini, obtained when certifying them for the Vanilla specification defined over a subgraph pivoted at the node for ‘Batman Begins’ movie. The samples reflect the certificates. Next, we identify and categorize prominent kinds of model responses. We frequently see the following failure modes — distracted and missed relation. In the former, the model gets deviated from the query by following the distractor context in its prompt, resulting in an incorrect answer. In the latter, the model skips some reasoning steps needed for the final correct answer. In cases of good reasoning, model accurately follows the query and gives the correct answer. Figure 11 presents examples of the aforementioned kinds of model responses for GPT-4o.
Figure 10: Qualitative analysis of samples used for certifying knowledge comprehension for Vanilla specifications on the Wikidata5m subgraph pivoted at the node for ‘Batman Begins’ movie. The context provided in the prompts is not shown for brevity. Wrong model responses are colored red and correct ones are colored green. The samples are consistent with our results, wherein Phi-3 (3B) has lower certification bounds than GPT-4o’s bounds, which are lower than those for Gemini-Pro.
Figure 11: GPT-4o responses showing good reasoning, failures due to distractors or skipped reasoning